# for Embodied Interactive Tasks

### v1

@@ Abstract
Recent advances in deep thinking models have demonstrated remarkable reasoning capabilities on mathematical and coding tasks. However, their effectiveness in embodied domains which require continuous interaction with environments through image action interleaved trajectories remains largely -unexplored. We present Embodied Reasoner, a model that extends o1 style reasoning to interactive embodied search tasks. Unlike mathematical reasoning that relies primarily on logical deduction, embodied scenarios demand spatial understanding, temporal reasoning, and ongoing self-reflection based on interaction history. To address these challenges, we synthesize 9.3k coherent Observation-Thought-Action trajectories containing 64k interactive images and 90k diverse thinking processes (analysis, spatial reasoning, reflection, planning, and verification). We develop a three stage training pipeline that progressively enhances the model’s capabilities through imitation learning, self exploration via rejection sampling, and self-correction through reflection tuning. The evaluation shows that our model significantly outperforms those advanced visual reasoning models, e.g., it exceeds OpenAI o1, o3-mini, and Claude-3.7 by +9%, 24%, and +13%. Analysis reveals our model exhibits fewer repeated searches and logical inconsistencies, with particular advantages in complex long horizon tasks. Real-world environments also show our superiority while exhibiting fewer repeated searches and logical inconsistency cases.
@@ 1 Introduction
Despite these advances, a critical question emerges: Can the o1-style reasoning paradigm be extended beyond these specialized domains to address more complex challenges that require embodied intelligence? Particularly, can these reasoning capabilities be effectively applied to embodied tasks demanding long-horizon planning and deliberate reasoning in interactive environments? This extension is non-trivial due to several fundamental challenges:
@@ 2.2 Instruction Synthesis
Task Templates with Constraints. We design multiple task templates for each task. Fig. 3 presents a transport task template: pickup {invisible A} and put in {B}, where A denotes a hidden object with pickupable attribute, e.g., keychain, and object B should contain containment properties, such as drawer or desk. It ensures synthesized instruction’s validity. Templates and constraints are in Appendix D.
@@ 4.1 Training Corpus
Our dataset spans 107 diverse indoor scenes, e.g., kitchens and living rooms, and covers 2,100 interactive objects (e.g., eggs, laptops) and 2,600 containers (e.g., refrigerators, drawers). All trajectories contain 64K a first-person perspective image from interaction and 8M thought tokens. The distribution of our dataset is presented in Appendix C.
@@ 5.3 Real-world Experiments
Fig. 7 illustrates an example: “Can you help me find the coffee and heat it up?” Our model rules out the countertop and dining table after two explorations (steps 1,2), ultimately locating the coffee (#7) in the cabinet and placing it in the microwave for heating (#11). However, we observe that OpenAI o3-mini fails to formulate a reasonable plan, heading to the microwave first instead of searching for the coffee. Besides, it frequently forgets to search and exhibits repetitive searching, aligning with our previous analysis. Please refer to Table B1 for detailed results.
@@ 6 Conclusions
Appendix A Related Works
A.1 Large Language Models Reasoning
Recent o1-style models [14, 30, 10, 62, 64, 39] have demonstrated powerful reasoning capabilities, significantly enhancing their ability on college-level benchmarks such as mathematics [35, 16]. Unlike previous efforts to scale up training data and model sizes, these systems involve generating long “thought” tokens during inference time, improving the performance of final answers [13, 27, 58, 60]. From chain-of-thought (CoT) prompting by humans [48, 22] to autonomous thinking derived by LLMs themselves, the reasoning abilities have become increasingly internalized and spontaneous. Besides, the long-thought generation process is quite similar to human thinking activities, with diverse thinking patterns, e.g., step-by-step solving, self-reflection, backtracking, and self-validation [32, 18]. These characteristics greatly enhance the complex problem-solving capabilities of LLMs. However, most efforts focus on text-based tasks, and the application of deep-thinking paradigms to embodied scenarios remains largely unexplored. This is the focus of our work.
A.2 Vision-Language Model Reasoning
A.3 Vision-Language-Action Model
Some research focuses on the integration of agents and language models. For instance, agents like PaLM-E[11] and other researches[25, 66, 56, 52, 65, 19, 28, 17, 33, 37, 6, 51] leverage the inherent knowledge of large language models (LLMs) and combine robotic data with general visual - language data for training. Another part of the research[9, 7] emphasizes the integration of language models (LLMs) and visual-anguage models (VLMs) into robotics. Take Chatgpt for robotics[42] as an example. The study first defines a high - level robotic function library, then creates prompts to guide ChatGPT to achieve its goals.
In addition, some research[50, 49, 63, 44] utilizes the capabilities of visual-language models (VLM) to assist traditional strategy training such as reinforcement learning and imitation learning, thus promoting the development of robotic manipulation.
Recently, a large amount of research[57, 8, 20, 43] has been dedicated to developing generalist robotic strategies, which are trained based on continuously expanding robotic learning datasets. However, most of the existing work directly predicts action sequences without enabling robots to have an autonomous thinking process and this is one of the key focuses of our work.
Appendix B Experiment Details
B.1 Real-World Experiments Result
To evaluate the generalization of our reasoning model, we design a real-world experiment about object searching, covering 30 tasks in three scenes. As shown in  Tab. B1, Embodied-Reasoner demonstrates notable capabilities in real-world environments. In terms of success rate, it outperforms OpenAI o1 by 6.7%, OpenAI o3-mini by 12.7%, and Qwen2.5-VL-72B-Instruct by 13.4%.
@@ B.1 Real-World Experiments Result
Table B1: The results of real-world experiments.
B.2 Repeat Exploration Rate
@@
Appendix C Dataset Details
C.1 Distribution of Task Instructions
@@
Figure C1: The distribution of the training dataset with 9,390 samples, including 4 task types and 10 sub-task types.
@@
We synthesize 9,390 unique task instructions along with their Observation-Thought-Action trajectories as the training set. The distribution of training tasks is shown in Fig. C1, encompassing 4 task types(Search, Manipulate, Transport and Composite) and 10 subtask types.
@@
For evaluation, we curate 809 test cases across 12 novel scenarios distinct from the training environments. The distribution of test tasks is shown in Fig. C2, covering 4 task types and 11 corresponding subtask types.
@@
Figure C2: The distribution of the test set with 809 tasks, including 4 task types and 11 sub-task types.
C.2 Distribution of Interaction Types
@@
In the training set, trajectories consist of eight types of interaction actions: navigate to, pickup, open, close, put in, observe, move forward, and toggle. As shown in the Fig. C3, the occurrence frequency of each interaction action across all trajectories is illustrated. Among them, the exploration action navigate to appears the most frequently, occurring over 29k times.
@@
In the test set, we manually design instructions and annotate the corresponding key actions and final states. The test tasks involve six types of interactions: navigate to, pickup, open, close, put in, and toggle. As seen in the Fig. C4, navigate to also appears significantly more frequently than other key actions.
@@
Figure C3: The distribution of the training set interactions, including 8 interaction types in trajectories: navigate to, pickup, open, close, put in, observe, move forward, and toggle.
@@
Figure C4: The distribution of the test set interactions, including 6 interaction types in key actions: navigate to, pickup, open, close, put in, and toggle.
C.3 Distribution of Task Length
@@
In the training set, each trajectory consists of an average of 7.2 interactions with the environment (e.g., navigate, pickup). For the four task types: Search, Manipulate, Transport, and Composite. Due to varying task complexity, the corresponding trajectory lengths also differ. As shown in the Fig. C5, Search tasks tend to have shorter trajectories, typically ranging from 1 to 9, while the more complex Composite tasks generate the longest trajectories, usually exceeding 8 and reaching beyond 23.
@@
Similarly, in the test set, as shown in the Fig. C6, the more complex Composite tasks also exhibit the longest key action sequences, usually exceeding 8 and reaching beyond 19.
@@
Figure C5: The quantity distribution of trajectory lengths in the training set and the corresponding task type composition, where Search Task is mainly within 1-9, Manipulate Task within 2-11, Transport Task within 3-14, and Composite Task above 8, extending beyond 23.
@@
Figure C6: The quantity distribution of key action lengths in the test set and the corresponding task type composition, where Search Task is mainly within 1-2, Manipulate Task within 2, 4-5, Transport Task within 4-7, and Composite Task above 8, extending beyond 19.
C.4 Distribution of Object Categories
@@
Our training dataset includes 107 indoor scenes with diverse functions and layouts (kitchens, living rooms, bedrooms, and bathrooms), featuring over 2,100 interactive objects (e.g., eggs, laptops) and 2,600 containers (e.g., refrigerators, drawers). Across the 9,390 unique task instructions, trajectories involve interactions with these objects and containers, with the top 32 most frequently explored and interacted object categories shown in Fig. C7. In the 12 distinct test set scenes, key actions correspond to different objects, with the top 32 most frequently involved containers and interactive objects illustrated in Fig. C8.
@@
Figure C7: The quantity distribution of the top 32 object types in the training dataset’s trajectories, with Others representing the remaining 62 categories, such as Bread, Book, DeskLamp, etc.
@@
Figure C8: The quantity distribution of the top 32 object types in the test set’s key actions, with Others representing the remaining 44 categories, such as Watch, Pencil, Cup, etc.
C.5 Description of Sub-task Types
@@
Appendix D Detailed Task Templates and Constraints
We design multiple task templates for each task. It ensures synthesized instruction’s validity. Templates and constraints are shown in Appendix D.
@@ Appendix D Detailed Task Templates and Constraints
Appendix E Detailed Prompts
@@
Appendix F Case Study
As shown in Figure F9,  Figure F10, and  Figure F11, we provide three cases. The first two illustrate different trajectory performances of Embodied-Reasoner and GPT-o1 for the same task, while the third presents Embodied-Reasoner’s performance in a real-world scenario.
F.1 Embodied-Reasoner’s Action Trajectories
As shown in Figure F9, it illustrates Embodied-Reasoner’s performance on the task: ”Could you please first place the Laptop on the Sofa, and then place the CellPhone on the Drawer?”
Figure F9: Trajectory Case for Embodied Reasoner
F.2 GPT-o1’s Action Trajectories
As shown in Figure F10, it illustrates GPT-o1’s performance on the task: ”Could you please first place the Laptop on the Sofa, and then place the CellPhone on the Drawer?”
@@ F.2 GPT-o1’s Action Trajectories
Figure F10: Trajectory Case for GPT-o1
F.3 Embodied-Reasoner in Real-World
As shown in Figure F11, it illustrates Embodied-Reasoner’s performance in the real world for the task: ”I’d appreciate it if you could leave the milk on the coffee table when possible”. It shows that Embodied-Reasoner can complete the task in real-world scenarios through step-by-step planning and reasoning.
Figure F11: Trajectory Case for Embodied Reasoner in Real World

### v2

@@ Abstract
Recent advances in deep thinking models have demonstrated remarkable reasoning capabilities on mathematical and coding tasks. However, their effectiveness in embodied domains which require continuous interaction with environments through image action interleaved trajectories remains largely -unexplored. We present Embodied Reasoner, a model that extends o1 style reasoning to interactive embodied search tasks. Unlike mathematical reasoning that relies primarily on logical deduction, embodied scenarios demand spatial understanding, temporal reasoning, and ongoing self-reflection based on interaction history. To address these challenges, we synthesize 9.3k coherent Observation-Thought-Action trajectories containing 64k interactive images and 90k diverse thinking processes (analysis, spatial reasoning, reflection, planning, and verification). We develop a three-stage training pipeline that progressively enhances the model’s capabilities through imitation learning, self-exploration via rejection sampling, and self-correction through reflection tuning. The evaluation shows that our model significantly outperforms those advanced visual reasoning models, e.g., it exceeds OpenAI o1, o3-mini, and Claude-3.7 by +9%, 24%, and +13%. Analysis reveals our model exhibits fewer repeated searches and logical inconsistencies, with particular advantages in complex long-horizon tasks. Real-world environments also show our superiority while exhibiting fewer repeated searches and logical inconsistency cases.
@@ 1 Introduction
Despite these advances, a critical question emerges: Can the o1-style reasoning paradigm be extended beyond these specialized domains to address more complex challenges that require embodied intelligence? Particularly, can these reasoning capabilities be effectively applied to embodied tasks demanding long-horizon planning and deliberate reasoning in interactive environments [44]? This extension is non-trivial due to several fundamental challenges:
@@ 2.2 Instruction Synthesis
Task Templates with Constraints. We design multiple task templates for each task. Fig. 3 presents a transport task template: pickup {invisible A} and put in {B}, where A denotes a hidden object with pickupable attribute, e.g., keychain, and object B should contain containment properties, such as drawer or desk. It ensures synthesized instruction’s validity. Templates and constraints are in Appendix C.
@@ 4.1 Training Corpus
Our dataset spans 107 diverse indoor scenes, e.g., kitchens and living rooms, and covers 2,100 interactive objects (e.g., eggs, laptops) and 2,600 containers (e.g., refrigerators, drawers). All trajectories contain 64K a first-person perspective image from interaction and 8M thought tokens. The distribution of our dataset is presented in Appendix B.
@@ 5.3 Real-world Experiments
Fig. 7 illustrates an example: “Can you help me find the coffee and heat it up?” Our model rules out the countertop and dining table after two explorations (steps 1,2), ultimately locating the coffee (#7) in the cabinet and placing it in the microwave for heating (#11). However, we observe that OpenAI o3-mini fails to formulate a reasonable plan, heading to the microwave first instead of searching for the coffee. Besides, it frequently forgets to search and exhibits repetitive searching, aligning with our previous analysis. Please refer to Table A1 for detailed results.
6 Related Works
6.1 Large Language Models Reasoning
Recent o1-style models [14, 35, 10, 72, 74, 47] have demonstrated powerful reasoning capabilities, significantly enhancing their ability on college-level benchmarks such as mathematics [40, 16]. Unlike previous efforts to scale up training data and model sizes, these systems involve generating long “thought” tokens during inference time, improving the performance of final answers [13, 32, 67, 69]. From chain-of-thought (CoT) prompting by humans [56, 25] to autonomous thinking derived by LLMs themselves [70], the reasoning abilities have become increasingly internalized and spontaneous. Besides, the long-thought generation process is quite similar to human thinking activities, with diverse thinking patterns, e.g., step-by-step solving, self-reflection, backtracking, and self-validation [37, 20]. These characteristics greatly enhance the complex problem-solving capabilities of LLMs. However, most efforts focus on text-based tasks, and the application of deep-thinking paradigms to embodied scenarios remains largely unexplored. This is the focus of our work.
6.2 Vision-Language Model Reasoning
6.3 Vision-Language-Action Model
Some researches integrate language models into embodied tasks [19, 28, 45, 41, 77]. For instance, agents like PaLM-E[11] and other researches[29, 76, 64, 60, 75, 21, 33, 18, 38, 43, 6, 59] leverage the inherent knowledge of large language models (LLMs) and combine robotic data with general visual-language data for training. Another part of the research[9, 7] emphasizes the integration of language models (LLMs) and visual-anguage models (VLMs) into robotics. Take Chatgpt for robotics[50] as an example. The study first defines a high - level robotic function library, then creates prompts to guide ChatGPT to achieve its goals.
In addition, some research[58, 57, 73, 52] utilizes the capabilities of visual-language models (VLM) to assist traditional strategy training such as reinforcement learning and imitation learning, thus promoting the development of robotic manipulation and navigation[17, 65, 23, 31]. Besides, a large amount of research[66, 8, 22, 51] has been dedicated to developing generalist robotic strategies, which are trained based on continuously expanding robotic learning datasets. However, most of the existing work directly predicts action sequences without enabling robots to have an autonomous thinking process and this is one of the key focuses of our work.
@@ 7 Conclusions
Appendix A Experiment Details
A.1 Real-World Experiments Result
To evaluate the generalization of our reasoning model, we design a real-world experiment about object searching, covering 30 tasks in three scenes. As shown in  Tab. A1, Embodied-Reasoner demonstrates notable capabilities in real-world environments. In terms of success rate, it outperforms OpenAI o1 by 6.7%, OpenAI o3-mini by 12.7%, and Qwen2.5-VL-72B-Instruct by 13.4%.
@@ A.1 Real-World Experiments Result
Table A1: The results of real-world experiments.
A.2 Repeat Exploration Rate
@@
Appendix B Dataset Details
B.1 Distribution of Task Instructions
@@
Figure B1: The distribution of the training dataset with 9,390 samples, including 4 task types and 10 sub-task types.
@@
We synthesize 9,390 unique task instructions along with their Observation-Thought-Action trajectories as the training set. The distribution of training tasks is shown in Fig. B1, encompassing 4 task types(Search, Manipulate, Transport and Composite) and 10 subtask types.
@@
For evaluation, we curate 809 test cases across 12 novel scenarios distinct from the training environments. The distribution of test tasks is shown in Fig. B2, covering 4 task types and 11 corresponding subtask types.
@@
Figure B2: The distribution of the test set with 809 tasks, including 4 task types and 11 sub-task types.
B.2 Distribution of Interaction Types
@@
In the training set, trajectories consist of eight types of interaction actions: navigate to, pickup, open, close, put in, observe, move forward, and toggle. As shown in the Fig. B3, the occurrence frequency of each interaction action across all trajectories is illustrated. Among them, the exploration action navigate to appears the most frequently, occurring over 29k times.
@@
In the test set, we manually design instructions and annotate the corresponding key actions and final states. The test tasks involve six types of interactions: navigate to, pickup, open, close, put in, and toggle. As seen in the Fig. B4, navigate to also appears significantly more frequently than other key actions.
@@
Figure B3: The distribution of the training set interactions, including 8 interaction types in trajectories: navigate to, pickup, open, close, put in, observe, move forward, and toggle.
@@
Figure B4: The distribution of the test set interactions, including 6 interaction types in key actions: navigate to, pickup, open, close, put in, and toggle.
B.3 Distribution of Task Length
@@
In the training set, each trajectory consists of an average of 7.2 interactions with the environment (e.g., navigate, pickup). For the four task types: Search, Manipulate, Transport, and Composite. Due to varying task complexity, the corresponding trajectory lengths also differ. As shown in the Fig. B5, Search tasks tend to have shorter trajectories, typically ranging from 1 to 9, while the more complex Composite tasks generate the longest trajectories, usually exceeding 8 and reaching beyond 23.
@@
Similarly, in the test set, as shown in the Fig. B6, the more complex Composite tasks also exhibit the longest key action sequences, usually exceeding 8 and reaching beyond 19.
@@
Figure B5: The quantity distribution of trajectory lengths in the training set and the corresponding task type composition, where Search Task is mainly within 1-9, Manipulate Task within 2-11, Transport Task within 3-14, and Composite Task above 8, extending beyond 23.
@@
Figure B6: The quantity distribution of key action lengths in the test set and the corresponding task type composition, where Search Task is mainly within 1-2, Manipulate Task within 2, 4-5, Transport Task within 4-7, and Composite Task above 8, extending beyond 19.
B.4 Distribution of Object Categories
@@
Our training dataset includes 107 indoor scenes with diverse functions and layouts (kitchens, living rooms, bedrooms, and bathrooms), featuring over 2,100 interactive objects (e.g., eggs, laptops) and 2,600 containers (e.g., refrigerators, drawers). Across the 9,390 unique task instructions, trajectories involve interactions with these objects and containers, with the top 32 most frequently explored and interacted object categories shown in Fig. B7. In the 12 distinct test set scenes, key actions correspond to different objects, with the top 32 most frequently involved containers and interactive objects illustrated in Fig. B8.
@@
Figure B7: The quantity distribution of the top 32 object types in the training dataset’s trajectories, with Others representing the remaining 62 categories, such as Bread, Book, DeskLamp, etc.
@@
Figure B8: The quantity distribution of the top 32 object types in the test set’s key actions, with Others representing the remaining 44 categories, such as Watch, Pencil, Cup, etc.
B.5 Description of Sub-task Types
@@
Appendix C Detailed Task Templates and Constraints
We design multiple task templates for each task. It ensures synthesized instruction’s validity. Templates and constraints are shown in Appendix C.
@@ Appendix C Detailed Task Templates and Constraints
Appendix D Detailed Prompts
@@
Appendix E Case Study
As shown in Figure E9,  Figure E10, and  Figure E11, we provide three cases. The first two illustrate different trajectory performances of Embodied-Reasoner and GPT-o1 for the same task, while the third presents Embodied-Reasoner’s performance in a real-world scenario.
E.1 Embodied-Reasoner’s Action Trajectories
As shown in Figure E9, it illustrates Embodied-Reasoner’s performance on the task: ”Could you please first place the Laptop on the Sofa, and then place the CellPhone on the Drawer?”
Figure E9: Trajectory Case for Embodied Reasoner
E.2 GPT-o1’s Action Trajectories
As shown in Figure E10, it illustrates GPT-o1’s performance on the task: ”Could you please first place the Laptop on the Sofa, and then place the CellPhone on the Drawer?”
@@ E.2 GPT-o1’s Action Trajectories
Figure E10: Trajectory Case for GPT-o1
E.3 Embodied-Reasoner in Real-World
As shown in Figure E11, it illustrates Embodied-Reasoner’s performance in the real world for the task: ”I’d appreciate it if you could leave the milk on the coffee table when possible”. It shows that Embodied-Reasoner can complete the task in real-world scenarios through step-by-step planning and reasoning.
Figure E11: Trajectory Case for Embodied Reasoner in Real World